# Misinterpreting Spin Precession as Orbital Eccentricity in Gravitational-Wave Signals

### v1

@@ Appendix B Settings for mismatch and PE studies
Our noise-weighed inner product between two signals $h_{1}$ and $h_{2}$ is the standard form for stationary Gaussian noise LIGOScientific:2025hdt; Veitch:2014wba; Thrane:2018qnx, and is used in both our mismatch study and in the GW likelihood used in PE. It is represented by (see also Thrane:2018qnx for discrete representations)
@@
Here, $S_{n}(f)$ is the PSD, and $f_{\mathrm{low}}$ and $f_{\mathrm{high}}$ are the lower and upper frequency cutoffs for the integral. For the mismatch studies we use the PSD for the Advanced LIGO design LIGO-T2200043; for the PE studies we inject the signal consistently into a network consisting of the Hanford, Livingston, and Virgo detectors, using their design PSDs LIGO-T2200043. For both the mismatch and PE studies we set $f_{\mathrm{low}}=20\,\mathrm{Hz}$. For the mismatch study we used $f_{\mathrm{high}}=2046$ Hz, and for PE we used $f_{\mathrm{high}}=1024$ Hz corresponding to a sample rate of $2048$ Hz; these cutoffs are expected to be sufficient to resolve higher modes during merger for the high-mass systems we study.
For the mismatch study we compare a precessing signal $h_{p}$ to eccentric signal $h_{e}$ in a single detector. Many of the intrinsic parameters are degenerate in this case Cutler:1994ys, but for concrete application we assume both signals to be located on the sky at spherical polar coordinates $\theta=\pi/3$ and $\varphi=\pi/6$ in the detector’s reference frame. For $h_{p}$ we set $\phi=0$, $\psi=0$, with arbitrary coalescence time $t=0$. These same parameters for $h_{e}$ are then set by the maximization within the mismatch.
@@
where $p(\boldsymbol{\theta}\mid d)$ is the posterior, $\mathcal{L}(d\mid\boldsymbol{\theta})$ is the likelihood, $\pi(\boldsymbol{\theta})$ is the prior, and $\mathcal{Z}$ is the evidence (marginal likelihood) which normalizes the probability density. We use the standard GW likelihood, e.g. Veitch:2014wba; Thrane:2018qnx; LIGOScientific:2025yae, which assumes a coherent signal across the detector network and which assumes independent, stationary, Gaussian noise in each detector as described by the PSD $S_{n}(f)$.

### v2

@@ Appendix B Settings for mismatch and PE studies
Our noise-weighed inner product between two signals $h_{1}$ and $h_{2}$ is the standard form for stationary Gaussian noise Abac and others (2025b); Veitch and others (2015); Thrane and Talbot (2019), and is used in both our mismatch study and in the GW likelihood used in PE. It is represented by (see also Thrane and Talbot (2019) for discrete representations)
@@
Here, $S_{n}(f)$ is the PSD, and $f_{\mathrm{low}}$ and $f_{\mathrm{high}}$ are the lower and upper frequency cutoffs for the integral. For the mismatch studies we use the PSD for the Advanced LIGO design 57; for the PE studies we inject the signal consistently into a network consisting of the Hanford, Livingston, and Virgo detectors, using their design PSDs 57. For both the mismatch and PE studies we set $f_{\mathrm{low}}=20\,\mathrm{Hz}$. For the mismatch study we used $f_{\mathrm{high}}=2046$ Hz, and for PE we used $f_{\mathrm{high}}=1024$ Hz corresponding to a sample rate of $2048$ Hz; these cutoffs are expected to be sufficient to resolve higher modes during merger for the high-mass systems we study.
For the mismatch study we compare a precessing signal $h_{p}$ to eccentric signal $h_{e}$ in a single detector. Many of the intrinsic parameters are degenerate in this case Cutler and Flanagan (1994), but for concrete application we assume both signals to be located on the sky at spherical polar coordinates $\theta=\pi/3$ and $\varphi=\pi/6$ in the detector’s reference frame. For $h_{p}$ we set $\phi=0$, $\psi=0$, with arbitrary coalescence time $t=0$. These same parameters for $h_{e}$ are then set by the maximization within the mismatch.
@@
where $p(\boldsymbol{\theta}\mid d)$ is the posterior, $\mathcal{L}(d\mid\boldsymbol{\theta})$ is the likelihood, $\pi(\boldsymbol{\theta})$ is the prior, and $\mathcal{Z}$ is the evidence (marginal likelihood) which normalizes the probability density. We use the standard GW likelihood, e.g. Veitch and others (2015); Thrane and Talbot (2019); Abac and others (2025c), which assumes a coherent signal across the detector network and which assumes independent, stationary, Gaussian noise in each detector as described by the PSD $S_{n}(f)$.